OpenStudy (anonymous):

The population of a region is growing exponentially. There were 20 million people in 1990 (t = 0) and 66 million in 2000. Find an expression for the population at any time t, in years. Use the general exponential function and remember to use exact values. P(t) = What population would you predict for the year 2010? What is the doubling time? (Round your answer to one decimal place.)

OpenStudy (ash2326):

Are you given the sample exponential function P(t)?

OpenStudy (anonymous):

no

OpenStudy (ash2326):

ok, exponential functions are of the form \[P(t)=A \times {10}^{Bt}\] we need to find A and B which are constants ! it's given that "There were 20 million people in 1990 (t = 0)" put t= 0 and P(t)= 20 million to find A Could you do that?

OpenStudy (anonymous):

So t=0 and b=??

OpenStudy (ash2326):

you need to find the value of A, once we get A then we'll find B

OpenStudy (anonymous):

20,000,000=A*10^B0

OpenStudy (ash2326):

yeah, so what would you get for A?

OpenStudy (anonymous):

please help me this is due in 10 mins :(

OpenStudy (ash2326):

@Brent0423 \[20000000=A 10^{B\times 0}\] \[20000000=A \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

8 mins :( im freaking out, can u please give me the answers then u can go back and explain

OpenStudy (ash2326):

8 minutes are enough, we'll solve it:) now you have A=20 million Population grows to 66 million in 2000 so t=2000-1990=10 years \[P(t)=A10^{BT}\] \[66million=10 million *10^{B\times 10}\] now find B, you'll need to use calculator

OpenStudy (anonymous):

theres 3 parts to this problem

OpenStudy (anonymous):

please just help me answer all 3 parts then we can go back, i will fail the assignment if i dont get this problem right

OpenStudy (anonymous):

4 mins :(

OpenStudy (ash2326):

you'll get \[66/20= 10^{B*10}\] \[B=\frac{\log 3.3}{10}\] it's just matter of 2 minutes, tell me the value of B

OpenStudy (anonymous):

.051

OpenStudy (ash2326):

now we need to find for 2010 t=2010-1990=20 so \[P(20)=20 million \times 10^{0.051\times 20}\] find population in 2010 from this

OpenStudy (ash2326):

B=0.081 I just checked

OpenStudy (ash2326):

no you were right sorry b=0.051 find the population

OpenStudy (ash2326):

@Brent0423 ??